OpenStudy (anonymous):

Help!! find x, 4x-6/y=15 if y=-2

OpenStudy (anonymous):

x=3/2y +15/4

OpenStudy (anonymous):

4x-6/(-2)=15 4x+3=15 -3 -3 4x=12 4 4 x=3